*Solving systems of Linear Equations in Three Variables by Elimination*
does anyone know how to solve x,y,z equations for x and explain?!
2x + y + z = 0
3x - y + z = 3
7x - 5y - 3z = 15
thanks!

Answer by stanbon(75887)

solve x,y,z equations for x and explain?!
2x + y + z = 0
3x - y + z = 3
7x - 5y - 3z = 15
-------
Setting up:
Subtract 1st row from 2nd and make the difference the 1st row:
x - 2y + 0 = 3
2x + y + z = 0
7x - 5y - 3z = 15
------------------------------
Subtract 2*1st row from 2nd row
Subtract 7*1st row from 3rd row
-------------------
x - 2y + 0 = 3
0 + 5y + z = -6
0 + 9y -3z = -6
-------
Solve equations 2 and 3 for y and z:
5y + z = -6
3y - z = -2
-----
Add and solve for "y":
8y = -8
y = -1
---
Solve for "z" using 3y-z = -2
3*-1 -z = -2
z = -1
-----
Solve for "x" using
2x + y + z = 0
2x + -1 + -1 = 0
x = 1
=========================
Solution:
x = 1
y = -1
z = -1
==============================
Cheers,
Stan H.
==============================